OpenStudy (1018):

Find the derivative: (x)(1-(4/x-3))

OpenStudy (1018):

(x)(4/(x-3)^2)+1(1-(4/x-3)) right?

OpenStudy (anonymous):

It should be -4x/(x-3)^2 + 1(1-(4/x-3))

OpenStudy (anonymous):

You just forgot to add the minus sign

OpenStudy (1018):

Ok thanks! Wait I'll try to solve again.

OpenStudy (1018):

I still got it without the minus sign.

OpenStudy (anonymous):

What is your actual question? I am not getting it..

OpenStudy (1018):

0-(-4) / (x-3)^2

OpenStudy (1018):

Find the derivative of the equation.

OpenStudy (anonymous):

\[x (1 - \frac{4}{x-3})\]

OpenStudy (1018):

I'll try to equate it to clarify.

OpenStudy (1018):

Haha, yes that's it.

OpenStudy (anonymous):

Multiply x inside first..

OpenStudy (anonymous):

\[1 - (\frac{(x-3) \cdot 4(1) - 4x \cdot 1}{(x-3)^2}) \implies \frac{(x-3)^2 - 4x + 12 + 4x}{(x-3)^2}\]

OpenStudy (anonymous):

You can simply cancel 4x with 4x there.. I hope I am getting your question properly.. :)

OpenStudy (anonymous):

LOOOOOOOOOOOOOOOOOOOOOOOOOOOOOL I'm so sorry. I thought it was 1 +4/(x-3) You are correct

OpenStudy (1018):

@hopelovelift lol that's alright, those +/- signs are annoying sometimes. Haha @waterineyes I'm a little lost. lol. I'll show you what I got so far, see if I'm in the right track

OpenStudy (1018):

\[(x)(\frac{ 4 }{ (x-3)^{2} }) + (1)(1 - \frac{ 4 }{ x-3 })\]

OpenStudy (anonymous):

What I got after simplification (x^2-6x+21)/(x-3)^2

OpenStudy (1018):

Oh man that's the correct answer. How'd you get that? Haha

OpenStudy (1018):

So I was wrong in the simplification. lol

OpenStudy (anonymous):

Its the same as waterineyes but a bit more simplified I'll show you, gimme a minute

OpenStudy (anonymous):

That is nothing but if you open the brackets for (x-3)^2

OpenStudy (anonymous):

And add 12 to it..

OpenStudy (anonymous):

\[(x-3)^2 = x^2 + 9 - 6x\] And add 12 to it.. :)

OpenStudy (anonymous):

 Simplify it more. I can't be stuffed drawing the rest XD All you are doing is making the denominator the same.